# Graph Neural Networks for Binary Programming

### v1

@@ Abstract
This paper investigates a link between Graph Neural Networks (GNNs) and Binary Programming (BP) problems, laying the groundwork for GNNs to approximate solutions for these computationally challenging problems. By analyzing the sensitivity of BP problems, we are able to frame the solution of BP problems as a heterophilic node classification task. We then propose Binary-Programming GNN (BPGNN), an architecture that integrates graph representation learning techniques with BP-aware features to approximate BP solutions efficiently. Additionally, we introduce a self-supervised data generation mechanism, to enable efficient and tractable training data acquisition even for large-scale BP problems. Experimental evaluations of BPGNN across diverse BP problem sizes showcase its superior performance compared to exhaustive search and heuristic approaches. Finally, we discuss open challenges in the under-explored field of BP problems with GNNs.
Binary Programming, Graph Neural Networks
@@ 1 Introduction
Solving Binary Programming (BP) Problems is of paramount importance in optimization and decision-making. These problems, characterized by binary decision variables, arise in numerous real-world scenarios, such as portfolio optimization (Black & Litterman, 1992), manufacturing and supply chain optimization (Geunes & Pardalos, 2005), telecommunications network optimization (Resende & Pardalos, 2008), resource allocation (Brown, 1984) and more. As the name suggests, Binary Programming (BP) is a type of mathematical problem where decision variables are restricted to take binary values (0 or 1). The discrete nature of BP makes these problems computationally challenging. Specifically, BP problems are known to be NP-complete (Karp, 2010). However, the solution of BP problems is crucial for optimizing the allocation of resources and improving operational efficiency. Contemporary solutions involve a range of optimization techniques, ranging from relaxation techniques and linear programming solvers (Nowak, 2005) to heuristic algorithms and heuristic-based approaches (Merz & Freisleben, 2002). Emerging technologies, such as quantum computing, show promise in addressing BP problems more efficiently (Glover et al., 2018), offering novel avenues for optimization in the future. However, they are not readily available yet, and their current setups are limited to solving problems with a small number of decision variables.
@@
Concurrently, in recent years, Graph Neural Networks (GNNs) have emerged as powerful machine learning models, and have been applied to various domains and applications, from social network analysis (Kipf & Welling, 2016; Defferrard et al., 2016; Veličković et al., 2018), bioinformatics (Jumper et al., 2021), to computer vision and graphics (Monti et al., 2017; Hanocka et al., 2019), and more.
@@
In this paper, we show that it is possible to model Binary Programming problems as a graph learning problem, namely node classification. Our findings lay the ground for leveraging the power of Graph Neural Networks in approximating the solutions of BP problems. Importantly, by characterizing the solution landscape of BP problems, we show that their solutions can be interpreted as heterophilic node classification. Therefore, we propose BPGNN, a Binary-Programming GNN, that builds on recent advances in heterophilic node classification with BP-aware features as part of its architecture. Another important aspect when using GNNs (or other data-driven techniques) to solve BP problems, is the scarcity of labeled data. While it is possible to use exhaustive search models to generate solutions of problems with a low number of decision variables, generating data for more than a dozen of decision variables becomes computationally intractable. We therefore propose a data generation procedure that stems from our theoretical analysis of the behavior of the solution landscape of BP problems. Our novel approach not only seeks to bridge the gap between traditional BP and cutting-edge neural graph representation learning techniques, but also holds the potential to offer more efficient and effective methodologies for addressing a wide range of computationally demanding optimization problems in diverse applications.
@@
Contributions: This paper makes several contributions on multiple fronts. (i) It establishes a novel connection that links between Graph Neural Networks and Binary Programming problems. This key observation serves as the foundation for exploring the utilization of GNNs in addressing and approximating solutions to computationally hard problems. (ii) We conduct a sensitivity analysis of the Binary Programming problems with respect to the input variables. This analysis sheds light on patterns and behavior of the solutions of BP learning problems, providing an understanding on how to address these problems using GNNs, by treating the solution of BP learning problems as heterophilic node classification. (iii) We propose an efficient self-supervised mechanism for generating training data, to address the NP-complete problem of generating data using exhaustive search otherwise. Finally, (iv) we evaluate BPGNN on varying sizes of BP problems and compare its performance with exhaustive search, heuristic, and other GNN approaches. Collectively, these contributions advance the understanding and utilization of GNNs in the realm of Binary Programming, paving the way for novel and efficient approaches to address this problem.
@@ 2 Problem Formulation
where ${\bf x}\in\mathbb{B}^{k}$ is a binary decision variables vector, ${\bf A}\in\mathbb{R}^{k\times k}$ is a real-valued matrix and ${\bf b}\in\mathbb{R}^{k}$ is a given real-valued vector, also sometimes called the observed vector. The QUBO problem was thoroughly studied in the literature, and we refer to (Beasley, 1998; Kochenberger et al., 2014; Glover et al., 2018) and reference within for theoretical as well as practical algorithms for its solution.
Obtaining exact solutions for QUBO problems is computationally challenging and limited, due to their discrete nature. As we discuss in Section 3, many heuristic-based methods have been proposed to solve this problem. However, despite such heuristics, it is still evident that in many cases, it is difficult to obtain a reasonable solution in polynomial time (Kochenberger et al., 2014), especially in high dimensions. In this work we focus on approximating the QUBO solution in a polynomial time for a subset of QUBO problems, assuming that the matrix ${\bf A}$ is sparse and fixed, but the observed vector ${\bf b}$ can vary. Such problems arise in virtually any field where QUBO problems are utilized (Black & Litterman, 1992; Geunes & Pardalos, 2005; Resende & Pardalos, 2008; Brown, 1984). For these problems, the matrix ${\bf A}$ represents the interaction between the variables, and the vector ${\bf b}$ is data that varies.
An important observation that we make is that QUBO can be interpreted as a graph problem. Let us consider a graph ${G=(V,E)}$ with a set of $k$ nodes, $V$, and a set of $m$ edges, $E$. The matrix ${\bf A}\in\mathbb{R}^{k\times k}$ from Equation 1 can be interpreted as the adjacency matrix of a graph ${G}$. The connectivity of ${G}$ is therefore represented by the sparsity pattern of ${\bf A}$, and the edge weights are represented by the matrix values. In the context of graphs and GNNs, we can treat ${\bf b}$ from Equation 1 as the input node features.
With this observation, we are motivated to explore the utilization of Graph Neural Networks (GNNs) to approximate the solution of QUBO problems. We show that through this approach, it is possible to efficiently obtain approximations to the solutions of the QUBO problem.
@@ 3 Related Work
3.1 Popular Binary Programming Solution Techniques
Solving QUBO problems has a wide and rich algorithmic treatment that ranges from exhaustive search methods, relaxation methods, differential equations based approaches, and grid search methods. We refer the reader to (Beasley, 1998) and more recently to (Goto et al., 2019) for a review of the algorithms. We now provide an overview of existing methods, focusing on representative methods that approximate a solution to the QUBO problem in Equation 1. We chose these techniques because they are publicly available and commonly used in literature Oshiyama & Ohzeki (2022).
Exhaustive Search. Overall, the QUBO problem has $2^{k}$ plausible solutions. While not always practical, the simplest approach is to use an exhaustive search. That is, test each of the possible solutions and choose the one that achieves the minimal value of Equation 1. This approach is viable when $k$ is small, but it becomes restrictive as $k$ grows. For instance, a problem of size $k=1024$ has $2^{1024}$ plausible solutions, making exhaustive search impractical.
Tabu Search (TS). Is a family of iterative algorithms that start from some initial guess ${\bf x}_{0}$, and defines a list, ${\cal T}$ of maximum length $\ell$, dubbed the Tabu list, that is initialized as with ${\bf x}_{0}$. The algorithm then defines a list of neighbors, ${\cal N}$ of ${\bf x}_{0}$ as perturbations in various policies (Beasley, 1998; Glover & Laguna, 1998; Glover et al., 2018). The simplest perturbation policy is defined as a vector $\tilde{{\bf x}}_{(i)}$ whose entries are identical to ${\bf x}_{0}$, with the exception of the $i$-th index. The neighbors list is then compared to the Tabu list and repeated entries are excluded from the neighbors list. That is, at iteration $t+1$ we can write, ${\cal N}_{t+1}={\cal N}_{t}\setminus{\cal T}_{t}$. The objective function is then evaluated on all the neighbors and the solution is updated according to
@@ 3.1 Popular Binary Programming Solution Techniques
The Tabu list is then updated, adding ${\bf x}_{t+1}$ to it and, possibly omitting the ${\bf x}_{t-\ell}$ term from it. The Tabu algorithm performs the described procedure for $T$ steps and returns the approximation vector with the lowest value obtained in the search. Note that each step in TS requires ${\cal O}(k)$ function evaluations. Although such an approach does not guarantee a global solution, it tends to yield low approximation error.
Simulating Adiabatic Bifurcations (SAB). A different approach that has been developed recently based on differential equations and molecular dynamics is SAB (Goto et al., 2019). In this approach, the decision variables in ${\bf x}$ are relaxed to be continuous, and a Hamiltonian function that has binary stationary points is minimized by simulating the dynamics to a steady state. This process is also known as annealing. In this approach, the cost of each step is a single matrix-vector product. However, as is typical with annealing methods, many steps are needed to reach convergence.
As the problem size, $k$, grows, it becomes difficult to obtain the exact solution with either TS or SAB (Oshiyama & Ohzeki, 2022). Furthermore, as we show in our experiments in Section 7, their computation time is high, making predictions in real time impractical using these approaches.
@@ 3.2 Graph Neural Networks
Graph Neural Networks (GNNs) were introduced in Scarselli et al. (2008) and became popular in recent years after advances in Defferrard et al. (2016); Kipf & Welling (2016). GNNs were shown to be effective in various applications and fields. A non-exhaustive list includes computer vision (Monti et al., 2017; Hanocka et al., 2019; Wang et al., 2018), social network analysis in semi-supervised settings (Kipf & Welling, 2016; Veličković et al., 2018), bioninformatics (Jumper et al., 2021; Corso et al., 2023). Essentially, GNNs allow the processing and learning from unstructured data described by graphs, to solve tasks such as node classification and graph classification, among other tasks. We refer the interested reader to Wu et al. (2020); Waikhom & Patgiri (2023) for a comprehensive review.
In this paper, we propose to use GNNs as a computational model to accelerate the solution approximation of BP problems, by viewing the problem as a heterophilic node classification task. We elaborate on this perspective in Section 4. Node classification problems are now a common application of GNNs (Kipf & Welling, 2016; Veličković et al., 2018; Pei et al., 2020; Chen et al., 2020; Chamberlain et al., 2021). It is common to categorize node classification problems into homophilic and heterophilic, or non-homophilic types. To define homophily and heterophily, we follow the definition in Pei et al. (2020), which measures the ground-truth label similarity of neighboring nodes. As shown in Gasteiger et al. (2019); Chamberlain et al. (2021), diffusion-based GNNs are useful for homophilic graphs, where neighboring nodes have similar labels. However, heterophilic node classification is deemed as a more challenging problem, and has been extensively studied in recent years. For example, some methods propose to learn both low and high pass filters (Chien et al., 2021; Luan et al., 2022; Eliasof et al., 2023b), while other methods focus on taking inspiration from neural ordinary differential equations (Haber et al., 2017; Chen et al., 2018) to incorporate operators that allow to model heterophilic behavior. Some recent examples are ACMP (Wang et al., 2022) and GREAD (Choi et al., 2023) that use reaction and diffusion terms to generate non-smooth patterns, as well as methods like CDE (Zhao et al., 2023) that utilize convection-diffusion dynamics, and ADR-GNN that uses advection-diffusion-reaction terms (Eliasof et al., 2023a). Other approaches include graph rewiring (Li et al., 2023), as well as learning directional message passing neural networks in Dir-GNN (Rossi et al., 2023). In the next section, we describe how a QUBO problem is transformed into a heterophilic node classification problem, and how to leverage the power of GNNs to solve such problems.
On another note, a recent line of works under the theme of Neural Algorithmic Reasoning (Veličković & Blundell, 2021; Ibarz et al., 2022; Bevilacqua et al., 2023) suggests to train a neural network to replicate the exact behavior of polynomial complexity algorithms. In this work, however, we seek to accelerate the computation of approximated solutions to the NP-complete QUBO problem.
@@ 4.1 Assumptions
Let ${\bf x}_{o}$ be the solution of the optimization problem in Equation 1. Given, ${\bf b}$, computing the exact solution requires solving an NP-complete optimization problem (Karp, 2010). Thus, it is common to compromise on an approximation solution, that can be obtained by classical solvers as described in Section 3. Still, those solvers require significant computational time and resources. We will now show that we are able to approximate the solutions using a neural approach, harnessing the power of GNNs.
Proposition 4.1.
@@
Let us assume that we have $n$ solved problems, for different observation vectors. Thus, we have observation-solution pairs $\{{\bf b}^{(j)},{\bf x}_{{o}}^{(j)}\}$, $j=0,\ldots,n-1$. Then, following Proposition 4.1, we suggest that we can approximate the function $q:\mathbb{R}^{k}\rightarrow{\mathbb{B}}^{k}$ via a neural network. Specifically, in the context of GNNs, the observed vector ${\bf b}$ and the solution ${\bf x}_{o}$ can be thought of as node features and labels, respectively. The graph connectivity,i.e., adjacency matrix, is defined by the sparsity of the matrix ${\bf A}$. Because our goal is to map every node to a binary value (0 or 1), the problem of learning the map $q$ can be formulated as a node classification problem, a common application of GNNs (Kipf & Welling, 2016).
4.2 Properties of the BP Problem
As stated in Proposition 4.1, we seek to utilize a node classification GNN to approximate BP solutions. Therefore, and in light of our discussion about homophilic and heterophilic datasets in Section 3.2, it is important to understand the properties of the solutions that we wish to approximate. We now analyze the sensitivity of the solution ${\bf x}_{o}$ with respect to the observed vector ${\bf b}$.
The Landscape of the Map $q(\cdot)$. To justify the ability to learn a map, $q:\mathbb{R}^{k}\rightarrow{\mathbb{B}}^{k}$, from the observed vector ${\bf b}$ to the BP solution, ${\bf x}_{o}$, we need to understand the behavior of the function $q({\bf b})$, and its sensitivity to changes in ${\bf b}$. While such analysis is not known to us for the general case, there are specific problems with physical interpretation that are thoroughly studied, as discussed in Example 1.

### v2

# Quadratic Binary Optimization with Graph Neural Networks

@@ Abstract
We investigate a link between Graph Neural Networks (GNNs) and Quadratic Unconstrained Binary Optimization (QUBO) problems, laying the groundwork for GNNs to approximate solutions for these computationally challenging tasks. By analyzing the sensitivity of QUBO formulations, we frame the solution of QUBO problems as a heterophilic node classification task. We then propose QUBO-GNN, an architecture that integrates graph representation learning techniques with QUBO-aware features to approximate solutions efficiently. Additionally, we introduce a self-supervised data generation mechanism to enable efficient and scalable training data acquisition even for large-scale QUBO instances. Experimental evaluations of QUBO-GNN across diverse QUBO problem sizes demonstrate its superior performance compared to exhaustive search and heuristic methods. Finally, we discuss open challenges in the emerging intersection between QUBO optimization and GNN-based learning.
[…]
1618
@@ 1 Introduction
Solving Binary Programming (BP) Problems is of paramount importance in optimization and decision-making. These problems, characterized by binary decision variables, arise in numerous real-world scenarios, such as portfolio optimization [3], manufacturing and supply chain optimization [17], telecommunications network optimization [43], resource allocation [4] and more. As the name suggests, Binary Programming (BP) is a type of mathematical problem where decision variables are restricted to take binary values (0 or 1). The discrete nature of BP makes these problems computationally challenging. Specifically, BP problems are known to be NP-complete [24]. However, the solution of BP problems is crucial for optimizing the allocation of resources and improving operational efficiency. Contemporary solutions involve a range of optimization techniques, ranging from relaxation techniques and linear programming solvers [37] to heuristic algorithms and heuristic-based approaches [32]. Emerging technologies, such as quantum computing, show promise in addressing BP problems more efficiently [19], offering novel avenues for optimization in the future. However, they are not readily available yet, and their current setups are limited to solving problems with a small number of decision variables. Other recent works proposed to combine reinforcement learning techniques [42] to learn a differentiable solver,
@@
Concurrently, in recent years, Graph Neural Networks (GNNs) have emerged as powerful machine learning models, and have been applied to various domains and applications, from social network analysis [27, 54], bioinformatics [22], to computer vision and graphics [34], and more.
@@
While early works suggested leveraging GNNs as an optimization framework to approximate the solution of relaxed versions of the BP problem [46], and Constrained Satisfaction Problems [29, 51], these approaches were restricted to a single instance of a problem and required the training from scratch of a network for a specific instance of a problem, or were done in an unsupervised manner [23]. Extending and departing from this approach, in this paper, we show that it is possible to model and characterize the family of solutions to BP problems as the solution of a graph learning task, namely, as heterophilic node classification. Our findings lay the ground for leveraging the power of Graph Neural Networks in approximating the solutions of BP problems in a supervised manner, allowing to generalize for new, unseen instances of BP problems. In this work, we focus on Quadratic Binary Optimization (QUBO) problems, which are a sub-family of BP, and we propose QUBO-GNN, a GNN architecture that builds on recent advances in heterophilic node classification with QUBO-aware features as part of its architecture. Another important aspect when using GNNs (or other data-driven techniques) to solve QUBO problems, is the scarcity of labeled data. While it is possible to use exhaustive search models to generate solutions of problems with a low number of decision variables, generating data for more than a dozen of decision variables becomes computationally intractable. We therefore propose a data generation procedure that stems from our theoretical analysis of the behavior of the solution landscape of QUBO problems. Our novel approach not only seeks to deepen the connection between QUBO and neural graph representation learning techniques, but also holds the potential to offer more efficient and effective methodologies for training a wide range of computationally demanding optimization problems in diverse applications.
@@
Contributions: This paper makes several contributions on multiple fronts. (i) It establishes a novel connection that links between Graph Neural Networks and Binary Programming problems. This key observation serves as the foundation for exploring the utilization of GNNs in addressing and approximating solutions to computationally hard problems. (ii) We conduct a sensitivity analysis of the Binary Programming problems with respect to the input variables. This analysis sheds light on patterns and behavior of the solutions of QUBO learning problems, providing an understanding of how to address these problems using GNNs, by treating the solution of QUBO learning problems as heterophilic node classification. (iii) We propose an efficient self-supervised mechanism for generating training data, to address the NP-complete problem of generating data using exhaustive search otherwise. Finally, (iv) we evaluate QUBO-GNN on varying sizes of QUBO problems and compare its performance with exhaustive search, heuristic, and other GNN approaches. Collectively, these contributions advance the understanding and utilization of GNNs in the realm of Binary Programming, paving the way for novel and efficient approaches to address this problem.
@@ 2 Problem Formulation
where ${\bf x}\in\mathbb{B}^{k}$ is a binary decision variables vector, ${\bf A}\in\mathbb{R}^{k\times k}$ is a real-valued matrix, and ${\bf b}\in\mathbb{R}^{k}$ is a given real-valued vector, also sometimes called the observed vector. The QUBO problem was thoroughly studied in the literature, and we refer to [2, 28, 19] and reference within for theoretical as well as practical algorithms for its solution.
Obtaining exact solutions for QUBO problems is computationally challenging, due to their discrete settings. As we discuss in Section 3, many heuristic-based methods have been proposed to solve this problem. However, despite such heuristics, it is still evident that in many cases, it is difficult to obtain a reasonable solution in polynomial time [28], especially in high dimensions. In this work, we focus on approximating the QUBO solution in a polynomial time for a subset of QUBO problems, assuming that the matrix $\bf A$} is sparse and fixed, but the observed vector $\bf b$} can vary. Such problems arise in virtually any field where QUBO problems are utilized [3, 17, 43, 4]. For these problems, the matrix $\bf A$} represents the interaction between the variables, and the vector $\bf b$} is data that varies.
An important observation that we make is that QUBO can be interpreted as a graph problem. Let us consider a graph G=(V,E){G=(V,E)} with a set of $k$ nodes, $V$, and a set of $m$ edges, $E$. The matrix ${\bf A}\in\mathbb{R}^{k\times k}$ from Equation˜1 can be interpreted as the adjacency matrix of a graph G{G}. The connectivity of G{G} is therefore represented by the sparsity pattern of $\bf A$}, and the edge weights are represented by the matrix values. In the context of graphs and GNNs, we can treat $\bf b$} from Equation˜1 as the input node features. With this observation, we are motivated to explore the utilization of Graph Neural Networks (GNNs) to approximate the solution of QUBO problems. We show that through this approach, it is possible to efficiently obtain approximations to the solutions of the QUBO problem.
@@ 3 Related Work
3.1 Popular QUBO Solution Techniques
Solving QUBO problems has a wide and rich algorithmic treatment that ranges from exhaustive search methods, relaxation methods, differential equations-based approaches, and grid search methods. We refer the reader to [2] and more recently to [20] for a review of the algorithms. We now provide an overview of existing methods, focusing on representative methods that approximate a solution to the QUBO problem in Equation˜1. We chose these techniques because they are publicly available and commonly used in literature [38].
Exhaustive Search.
Overall, the QUBO problem has $2^{k}$ plausible solutions. While not always practical, the simplest approach is to use an exhaustive search. That is, test each of the possible solutions and choose the one that achieves the minimal value of Equation˜1. This approach is viable when $k$ is small, but it becomes restrictive as $k$ grows. For instance, a problem of size $k=1024$ has $2^{1024}$ plausible solutions, making exhaustive search impractical.
Tabu Search (TS).
Is a family of iterative algorithms that start from some initial guess ${\bf x}_{0}$, and defines a list, $\cal T$} of maximum length $\ell$, dubbed the Tabu list, that is initialized as with ${\bf x}_{0}$. The algorithm then defines a list of neighbors, $\cal N$} of ${\bf x}_{0}$ as perturbations in various policies [2, 18, 19]. The simplest perturbation policy is defined as a vector $\tilde{{\bf x}}_{(i)}$ whose entries are identical to ${\bf x}_{0}$, with the exception of the $i$-th index. The neighbors list is then compared to the Tabu list, and repeated entries are excluded from the neighbors list. That is, at iteration $t+1$ we can write ${\cal N}_{t+1}={\cal N}_{t}\setminus{\cal T}_{t}$. The objective function is then evaluated on all the neighbors and the solution is updated according to
@@ 3.1 Popular QUBO Solution Techniques
The Tabu list is then updated, adding ${\bf x}_{t+1}$ to it and, possibly omitting the ${\bf x}_{t-\ell}$ term from it. The Tabu algorithm performs the described procedure for $T$ steps and returns the approximation vector with the lowest value obtained in the search. Note that each step in TS requires ${\cal O}(k)$ function evaluations. While this approach does not guarantee a global solution, it tends to yield low approximation error.
Simulating Adiabatic Bifurcations (SAB).
A different approach that has been developed recently based on differential equations and molecular dynamics is SAB [20]. In this approach, the decision variables in $\bf x$} are relaxed to be continuous, and a Hamiltonian function that has binary stationary points is minimized by simulating the dynamics to a steady state. This process is also known as annealing. In this approach, the cost of each step is a single matrix-vector product. However, as is typical with annealing methods, many steps are needed to reach convergence. As the problem size, $k$, grows, it becomes difficult to obtain the exact solution with either TS or SAB [38]. Furthermore, as we show in our experiments in Section 7, their computation time is high, making predictions in real-time impractical using these approaches.
@@ 3.2 Graph Neural Networks
Graph Neural Networks (GNNs) were introduced in Scarselli et al. [45] and became popular in [27]. GNNs were shown to be effective in various applications and fields. A non-exhaustive list includes computer vision [56], social network analysis in semi-supervised settings [27, 54], bioninformatics [22]. Essentially, GNNs allow the processing and learning from unstructured data described by graphs, to solve tasks such as node classification and graph classification, among other tasks. We refer the interested reader to Waikhom and Patgiri [55] for a comprehensive review. Therefore, it was proposed in Schuetz et al. [46] to leverage GNNs as an optimization framework for the solution of relaxations of instances of QUBO problems. In these approaches, the idea is to train a GNN from scratch for every pair $({\bf A},{\bf b})$.
Uniquely, we propose to use GNNs as a computational model to accelerate and generalize the solution approximation of QUBO problems, by viewing the problem as a heterophilic node classification task. We elaborate on this perspective in Section 4. Node classification problems are now a common application of GNNs [27, 54, 40, 5]. It is common to categorize node classification problems into homophilic and heterophilic, or non-homophilic types. To define homophily and heterophily, we follow the definition in Pei et al. [40], which measures the ground-truth label similarity of neighboring nodes. As shown in Chamberlain et al. [5], diffusion-based GNNs are useful for homophilic graphs, where neighboring nodes have similar labels. However, heterophilic node classification is deemed as a more challenging problem, and has been extensively studied in recent years. For example, some methods propose to learn both low and high pass filters [13], while other methods focus on taking inspiration from neural ordinary differential equations [7] to incorporate operators that allow to model heterophilic behavior. A recent example is GREAD [8] that use reaction and diffusion terms to generate non-smooth patterns, as well as methods like CDE [58] that utilize convection-diffusion dynamics. Other approaches consider learning directional message passing neural networks, as in Dir-GNN [44]. In the next section, we describe how a QUBO problem is transformed into a heterophilic node classification problem, and how to leverage the power of GNNs to solve such problems.
On another note, a recent line of works under the theme of Neural Algorithmic Reasoning [53] suggests to train a neural network to replicate the exact behavior of polynomial complexity algorithms. In this work, however, we seek to accelerate the computation of approximated solutions to the NP-complete QUBO problem.
@@ 4.1 Assumptions
Let ${\bf x}_{o}$ be the solution of the optimization problem in Equation˜1. Given, $\bf b$}, computing the exact solution requires solving an NP-complete optimization problem [24]. Thus, it is common to compromise on an approximation solution, that can be obtained by classical solvers as described in Section 3. Still, those solvers require significant computational time and resources. We will now show that we are able to approximate the solutions using a neural approach, harnessing the power of GNNs.
Proposition 1.
@@
Let us assume that we have $n$ solved problems, for different observation vectors. Thus, we have observation-solution pairs $\{{\bf b}^{(j)},{\bf x}_{{o}}^{(j)}\}$, $j=0,\ldots,n-1$. Then, following Proposition 1, we suggest that we can approximate the function $q:\mathbb{R}^{k}\rightarrow{\mathbb{B}}^{k}$ via a neural network. If that was not the case, then by theory and experimental findings, learning $q$ would be a highly challenging task [47, 41, 48]. Specifically, in the context of GNNs, the observed vector $\bf b$} and the solution ${\bf x}_{o}$ can be thought of as node features and labels, respectively. The graph connectivity, i.e., adjacency matrix, is defined by the sparsity of the matrix $\bf A$}. Because our goal is to map every node to a binary value (0 or 1), the problem of learning the map $q$ can be formulated as a node classification problem, a common application of GNNs [27].
4.2 Properties of the QUBO Problem
As stated in Proposition 1, we seek to utilize a node classification GNN to approximate QUBO solutions. Therefore, and in light of our discussion about homophilic and heterophilic datasets in Section 3.2, it is important to understand the properties of the solutions that we wish to approximate. We now analyze the sensitivity of the solution ${\bf x}_{o}$ with respect to the observed vector $\bf b$}.
The Landscape of the Map $q(\cdot)$.
To justify the ability to learn a map, $q:\mathbb{R}^{k}\rightarrow{\mathbb{B}}^{k}$, from the observed vector $\bf b$} to the QUBO solution, ${\bf x}_{o}$, we need to understand the behavior of the function $q({\bf b})$, and its sensitivity to changes in $\bf b$}. While such analysis is not known to us for the general case, there are specific problems with physical interpretation, as in Example˜1.